# Towards Invisible Backdoor Attack on Text-to-Image Diffusion Model

### v1

@@ Abstract
Backdoor attacks targeting text-to-image diffusion models have advanced rapidly, enabling attackers to implant malicious triggers into these models to manipulate their outputs. However, current backdoor samples often exhibit two key abnormalities compared to benign samples: 1) Semantic Consistency, where backdoor prompts tend to generate images with similar semantic content even with significant textual variations to the prompts; 2) Attention Consistency, where the trigger induces consistent structural responses in the cross-attention maps. These consistencies leave detectable traces for defenders, making backdoors easier to identify. To enhance the stealthiness of backdoor samples, we propose a novel Invisible Backdoor Attack (IBA) by explicitly mitigating these consistencies. Specifically, our approach leverages syntactic structures as backdoor triggers to amplify the sensitivity to textual variations, effectively breaking down the semantic consistency. Besides, a regularization method based on Kernel Maximum Mean Discrepancy (KMMD) is proposed to align the distribution of cross-attention responses between backdoor and benign samples, thereby disrupting attention consistency. Extensive experiments demonstrate that our IBA achieves a 97.5% attack success rate while exhibiting stronger resistance to defenses, with an average of over 98% backdoor samples bypassing three state-of-the-art detection mechanisms. The code is available at https://github.com/Robin-WZQ/IBA.
1 Introduction
In recent years, Text-to-Image (T2I) diffusion models have made remarkable progress [18, 40, 9, 17, 34, 35, 47], demonstrating the favorable ability of generating high-quality images through a text-guided denoising process. They have been successfully applied to various downstream tasks, including image generation [31, 11], image editing [16, 6], and video generation [27].
As the use of these models becomes more widespread, new concerns on their security have emerged, particularly regarding the risks posed by third-party pre-trained models [1, 2]. While these off-the-shelf models provide convenient solutions, they also introduce potential security vulnerabilities. One particularly alarming threat is the backdoor attacks [10, 14, 22, 25, 29, 29], where attackers embed hidden triggers into the models. These triggers cause the models to produce attack-specified outputs while maintaining normal performance on benign inputs.
@@ 1 Introduction
Although backdoor attacks on T2I diffusion models achieve high attack success rates, they still exhibit detectable abnormalities, making them easier to identify. Two common abnormalities are: 1) Semantic Consistency, where backdoor prompts generate images with similar semantic content even when words are added to or removed from the prompts [15]; 2) Attention Consistency, where the triggers induce the consistent structural responses in the cross-attention map [16, 13, 44], as illustrated in the middle row of Fig. 3. These characteristics provide defenders with effective cues for detecting backdoor samples [44, 15], reducing the stealthiness of backdoor attacks.
In this work, we propose a novel Invisible Backdoor Attack (IBA) method toward improving the stealthiness of backdoor samples, involving dual-modal optimization of textual features and diffusion features on the text encoder. Firstly, inspired by [32], we introduce a syntactic structure-based backdoor attack, leveraging specific syntactic patterns in natural language as triggers. In contrast to utilizing the specific tokens as trigger [41, 20, 8, 42, 46], rigorous syntactic triggers enhance the sensitivity of backdoor samples to textual perturbations, thereby disrupting semantic consistency. Furthermore, we introduce a loss term based on Kernel Mean Matching Distance (KMMD) [39]. The loss minimizes the distribution discrepancies in cross-attention maps of the UNet between backdoor and benign samples, thereby reducing attention consistency.
To evaluate the effectiveness of IBA, we conduct extensive experiments focusing on both attack success rate and detection resistance. Specifically, we assess IBA against four state-of-the-art backdoor defense algorithms, including T2IShield-FTT [44], T2IShield-CDA [44], UFID [15] and textual perturbation [7]. The experimental results show that IBA achieves comparable attack performance to existing backdoor attack methods [41, 20, 8, 42, 46], with an average attack success rate of 97.5%. Besides, an average of 98% backdoor samples generated by IBA bypass detection mechanisms. Extensive experiments are also conducted for exploring the effectiveness of IBA. We highlight our main contributions as follows:
  • We propose a novel method towards Invisible Backdoor Attack (IBA) on Text-to-Image diffusion models, which leveraging dual-modal features to jointly optimize the injection, significantly improving stealthiness while maintaining high attack success rate.
@@
2 Related Works
2.1 Text-to-Image Diffusion Model
@@
2.2 Backdoor Attacks on T2I Diffusion Models
@@
Backdoor attacks have been extensively studied on various tasks, particularly on the image classification [10, 14, 22, 25, 29]. These attacks aim to implant triggers into a victim model, enabling attackers to manipulate its output while maintaining performance for benign inputs. Recently, several works have been explored backdoor attacks on T2I diffusion models [41, 20, 8, 42, 46, 48]. Struppek et al. introduce Rickrolling the Artist [41], which involves embedding a visually similar characters by align the features between poisoned and target prompts in the text embedding space. Huang et al. [20] and Wu et al. [46] utilize personalization techniques [20] to implant word combinations into the model. Zhai et al. introduce BadT2I [48]. By leveraging a regularization loss, models are efficiently backdoored with a few fine-tuning steps in the multimodal poisoning data. Chou et al. [8] introduce VillanDiffusion, a method that integrates trigger implantation into LoRA [19]. Unlike previous works relying on poisoned data, Wang et al. [43] propose EvilEdit, a training-free and data-free approach that implants backdoors by modifying the weights.
2.3 Backdoor Defense on T2I Diffusion Models
@@
In response to the increasing security threat posed by backdoor attacks, several defense mechanisms have been proposed. Wang et al. [44] propose a comprehensive framework named T2Ishield that detects, localizes, and mitigates backdoor samples by identifying the “Assimilation Phenomenon”, i.e., a consistent structural response in attention maps. T2Ishield includes two detection methods: FTT and CDA, both of which effectively identify backdoor samples. Chew et al. [7] further proposed a defense method based on text perturbation, where applying character-level and word-level perturbations to backdoor samples. Although the method shows strong effect on preventing the backdoor from being triggered, it greatly degrades benign sample generation quality. Furthermore, Guan et al. [15] introduce UFID, a novel defense method that leverages output diversity as a metric to differentiate backdoor samples from benign ones. UFID is based on the observation that backdoor samples are less sensitive to textual variations. In this work, we propose an invisible backdoor attack to improve stealthiness while still maintaining the strong backdoor attack performance.
Figure 1: The overview of Invisible Backdoor Attack (IBA). (a) Data Generation: given a syntactic template, we leverage GPT-4o [3] to generate human-readable backdoor samples $\hat{x}_{i}$ that match the template. Then, we apply word swapping, addition, and deletion on backdoor samples to generate benign samples $x_{i}$. (b) Backdoor Injection: we inject the backdoor via jointly optimizing three losses: $\mathcal{L}_{Benign}$ for minimizing the embedding distance of benign samples between the clean text encoder $F_{\theta}(\cdot)$ and the backdoor text encoder $\hat{F}_{\theta}(\cdot)$, $\mathcal{L}_{Backdoor}$ for aligning the embedding of backdoor samples to target sample, $\mathcal{L}_{KMMDR}$ for providing the gradient information from the U-Net [36] for better aligning cross-attention responses of the diffusion process between backdoor and benign samples.
3 Methods
@@
In this paper, we propose a novel Invisible Backdoor Attack (IBA) framework, which significantly enhances the stealthiness of backdoor samples. We start with an overview of our IBA, then describe the process of generating backdoor and benign samples, and finally introduce three optimization objectives used for backdoor injection.
3.1 Overview of Our Methods
@@
Following prior works [41, 8, 20, 44], we consider Text-to-Image (T2I) diffusion models [35] as the target models for backdoor attacks. In this scenario, the attacker, who has access to the model parameters, injects backdoors into the model and subsequently releases the modified model on the third-party platforms [1].
@@
The overview of our IBA is shown in Fig. 1. Given a T2I diffusion model, the goal is to inject the backdoors into the text encoder $F_{\theta}(\cdot)$ of T2I diffusion models (i.e., CLIP [33]). The backdoor model $\hat{F}_{\theta}(\cdot)$ is expected to output embedding that guides the T2I diffusion model to generate attacker-specified images while mitigating abnormal consistencies. Specifically, IBA involves two stages, i.e., data generation and backdoor injection. During the data generation, we first construct a dataset of backdoor samples $\mathbb{\hat{D}}=\{\hat{x}_{i},\hat{y}\}$, where $\hat{x}_{i}$ represents the prompt with the specific syntactic structure and $\hat{y}$ denotes the attacker-specified target prompt. Benign samples $\mathbb{D}=\{x_{i}\}$ are then generated by applying word-order swapping, words addition, and words deletion transformations to the backdoor samples $\hat{x}_{i}$. For the backdoor injection, the clean model $F_{\theta}(\cdot)$ is fine-tuned on both $\mathbb{\hat{D}}$ and $\mathbb{D}$ to obtain the backdoor model $\hat{F}_{\theta}(\cdot)$. Besides, a regularization loss $\mathcal{L}_{KMMDR}$ based on Kernel Mean Matching Distance (KMMD) is proposed for mitigating the abnormal attention consistency. The overall objective of our method is formulated as follows:
@@
3.2 Samples Generation w & w/o Syntactic Trigger
@@
Trigger Syntactic Templates Selection. Based on previous researches [41, 32], we select rare syntactic templates to generate backdoor samples that effectively distinguish from benign samples. Specifically, we analyze DiffusionDB dataset [45], which contains high-quality prompts from real users. We then leverage the Stanford Parser [28] to parse each prompt and calculate the frequency of various syntactic structures. We choose structures that occur infrequently, such as (DET) (NOUN) (ADP) (DET) (NOUN) (VERB) (ADP) (NOUN), to serve as our templates.
Figure 2: The visualization of cross-attention maps during image generation. Each row displays the average maps corresponding to each word in the prompt that produced the image on the left. (Top) A benign sample. (Middle) A backdoor sample trained by Rickrolling [41]. (Bottom) A backdoor sample with the syntactic trigger trained by IBA.
@@
Backdoor Samples Generation. With the selected syntactic templates, we aim to generate a set of backdoor samples $\mathbb{\hat{D}}$ that conform to these syntactic structures. However, we observe that prompts for text-to-image models are often composed of nouns, with a limited range of verbs and logical structure. This makes it challenging to generate human-readable prompts through rule-based methods [32]. To address this, we utilize GPT-4o [3] to assist in the generation process. We instruct the model to generate sentences with low perplexity that are human-readable, cover varying scenarios, and incorporate varied vocabulary. Finally, we manually filter out incorrectly generated samples. More generated prompts can be found in Sec. A.7.
Benign Samples Generation. Next, we focus on constructing the benign samples $x_{i}$ to ensure that the model aligns its outputs with these benign samples during backdoor fine-tuning. Instead of directly sampling prompts directly from DiffusionDB [45], we generate benign samples by applying operations such as word-order swapping, words addition, and words deletion to the backdoor samples $\hat{x_{i}}$. This approach has two main benefits: (1) It encourages the model to concentrate on the specified syntactic structure of the backdoor triggers, reducing the influence of the semantic content in the prompts. (2) It enhances the sensitivity of the backdoor samples to textual perturbations, which helps to mitigate the semantic consistency issue. We provide further experiments to support the benefits of operations in Sec. A.4.
3.3 Backdoor Injection
During the backdoor injection process, we utilize a teacher-student paradigm to implant the backdoor, where both models are initialized using a pre-trained CLIP [33]. The student model serves as the poisoned model, while the teacher model is used to align the feature embeddings of samples from $\mathbb{\hat{D}}$ and $\mathbb{D}$ with the student model. It is noted that IBA also involves dual-modal optimization, leveraging gradient from the UNet [36] during the diffusion generation process to jointly optimize the backdoor injection. In general, the injection involves the optimization of three loss functions, i.e., Benign Loss, Backdoor Loss and KMMDR Loss.
@@ 3.3 Backdoor Injection
where $x,y\in[1,D]$, and $\bar{M}$ is the mean of the attention maps across tokens. Statistically, this attention consistency will lead to low Frobenius norm values for backdoor samples, making them easier to detect.
@@
Finally, the optimization objective of our IBA is:
@@
where $\gamma$ and $\lambda$ are parameters that balance the contributions of the Benign Loss, Backdoor Loss and KMMDR Loss.
4 Experiments
Figure 4: Qualitative results of IBA. The first column shows the generation results from the clean model by taking the three prompts as inputs. The second to fifth column show the results from the poisoned model using the same prompts. Each poisoned model is injected with a different backdoor target. We provide more examples in Sec. A.7.

### v2

# Trigger without Trace: Towards Stealthy Backdoor Attack on Text-to-Image Diffusion Models

@@ Abstract
Backdoor attacks targeting text-to-image diffusion models have advanced rapidly. However, current backdoor samples often exhibit two key abnormalities compared to benign samples: 1) Semantic Consistency, where backdoor prompts tend to generate images with similar semantic content even with significant textual variations to the prompts; 2) Attention Consistency, where the trigger induces consistent structural responses in the cross-attention maps. These consistencies leave detectable traces for defenders, making backdoors easier to identify. In this paper, toward stealthy backdoor samples, we propose Trigger without Trace (TwT) by explicitly mitigating these consistencies. Specifically, our approach leverages syntactic structures as backdoor triggers to amplify the sensitivity to textual variations, effectively breaking down the semantic consistency. Besides, a regularization method based on Kernel Maximum Mean Discrepancy (KMMD) is proposed to align the distribution of cross-attention responses between backdoor and benign samples, thereby disrupting attention consistency. Extensive experiments demonstrate that our method achieves a 97.5% attack success rate while exhibiting stronger resistance to defenses. It achieves an average of over 98% backdoor samples bypassing three state-of-the-art detection mechanisms, revealing the vulnerabilities of current backdoor defense methods. The code is available at https://github.com/Robin-WZQ/TwT.
Index Terms:
Backdoor Attack, Text-to-Image Diffusion Models, Syntactic Trigger.
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Recent years have witnessed significant advances in Text-to-Image (T2I) diffusion models [1, 2, 3, 4, 5, 6, 7, 8, 9]. These models are designed to synthesize high-fidelity images guided by natural language prompts, leveraging a denoising process conditioned on text embeddings. Their remarkable generative capabilities have facilitated a wide range of downstream applications, including image generation [10, 11], image editing [12, 13], and video generation [14].
As the use of these models becomes more widespread, new concerns on their security have emerged, particularly regarding the risks posed by third-party pre-trained models [15, 16]. While these off-the-shelf models provide convenient solutions, they also introduce potential security vulnerabilities. One particularly alarming threat is the backdoor attacks [17, 18, 19, 20, 21, 22, 23, 24], where attackers embed hidden triggers into the models. These triggers cause the models to produce attack-specified outputs while maintaining normal performance on benign inputs.
@@ I Introduction
Although backdoor attacks on T2I diffusion models achieve high attack success rates, they still exhibit detectable abnormalities. Two common abnormalities are: ❶ Semantic Consistency, where backdoor prompts generate images with similar semantic content even when words are added to or removed from the prompts [32]; ❷ Attention Consistency, where the triggers induce the consistent structural responses in the cross-attention map [12, 33, 34], as illustrated in the middle row of Fig. 3. These characteristics leave detectable traces that can be used to identify backdoor samples [34, 32], thereby reducing the stealthiness of backdoor attacks.
In this work, we propose Trigger without Trace (TwT) to improve the stealthiness of backdoor samples, revealing the vulnerability of current defense methods. The method involves dual-modal optimization of textual features and diffusion features on the text encoder. Firstly, inspired by [35], we introduce a syntactic structure-based backdoor attack, leveraging specific syntactic patterns in natural language as triggers. In contrast to utilizing the specific tokens as trigger [25, 26, 27, 24, 28], rigorous syntactic triggers enhance the sensitivity of backdoor samples to textual perturbations, thereby disrupting semantic consistency. Furthermore, we introduce a loss term based on Kernel Mean Matching Distance (KMMD) [36]. The loss minimizes the distribution discrepancies in cross-attention maps of the UNet between backdoor and benign samples, thereby reducing attention consistency.
To evaluate the effectiveness of our method, we conduct extensive experiments focusing on both attack success rate and detection resistance. Specifically, we assess our method against four state-of-the-art backdoor defense algorithms, including T2IShield-FTT [34], T2IShield-CDA [34], UFID [32] and textual perturbation [37]. The experimental results show that our method achieves comparable attack performance to existing backdoor attack methods [25, 26, 27, 24, 28], with an average attack success rate of 97.5%. Besides, an average of 98% backdoor samples generated by our method bypass detection mechanisms, i.e., T2IShield and UFID. We hope our work can help identify vulnerabilities in current backdoor defense methods and raise awareness of potential backdoor risks within the community.
To conclude, our main contributions are as follows:
  • We propose Trigger without Trace (TwT) on Text-to-Image diffusion models, which leveraging dual-modal features to jointly optimize the injection, significantly improving stealthiness while maintaining high attack success rate.
@@
II Related Works
In this section, we first discuss the development of text-to-image diffusion models. Then, we review the landscape of both backdoor attack and defense methods on T2I diffusion models, respectively.
II-A Text-to-Image Diffusion Model
@@
II-B Backdoor Attacks on T2I Diffusion Models
@@
Backdoor attacks have been extensively studied on various tasks, particularly on the classification tasks[17, 18, 19, 20, 21]. These attacks aim to implant triggers into a victim model, enabling attackers to misclassify to a specific label while maintaining performance for benign inputs. Recently, several works have been explored backdoor attacks on T2I diffusion models [25, 26, 27, 24, 28, 29]. Struppek et al. introduce Rickrolling the Artist [25], which involves embedding a visually similar characters by align the features between poisoned and target prompts in the text embedding space. Huang et al. [26] and Wu et al. [28] utilize personalization techniques [26] to implant word combinations into the model. Zhai et al. introduce BadT2I [29]. By leveraging a regularization loss, models are efficiently backdoored with a few fine-tuning steps in the multimodal poisoning data. Chou et al. [27] introduce VillanDiffusion, a method that integrates trigger implantation into LoRA [31]. Unlike previous works relying on poisoned data, Wang et al. [30] propose EvilEdit, a training-free and data-free approach that implants backdoors by modifying the weights.
II-C Backdoor Defense on T2I Diffusion Models
@@
In response to the increasing security threat posed by backdoor attacks, several defense mechanisms have been proposed. Wang et al. [34] propose a comprehensive framework named T2IShield that detects, localizes, and mitigates backdoor samples by identifying the “Assimilation Phenomenon”, i.e., a consistent structural response in attention maps. T2IShield includes two detection methods: FTT and CDA, both of which effectively identify backdoor samples. Chew et al. [37] further proposed a defense method based on text perturbation, where applying character-level and word-level perturbations to backdoor samples. Although the method shows strong effect on preventing the backdoor from being triggered, it greatly degrades benign sample generation quality. Furthermore, Guan et al. [32] introduce UFID, a novel defense method that leverages output diversity as a metric to differentiate backdoor samples from benign ones. In this work, we propose Trigger without Trace to improve stealthiness while still maintaining the strong backdoor attack performance.
III Methods
@@
In this paper, we propose a novel backdoor attack framework, which significantly enhances the stealthiness of backdoor samples. We start with an overview of our method in Section III-A, then describe the process of generating backdoor and benign samples in Section 3, and finally introduce three optimization objectives used for backdoor injection in Section III-C.
Figure 1: The overview of our method. (a) Data Generation: given a syntactic template $s$, we leverage a off-the-shelf LLM [43] $\mathcal{G}$ to generate human-readable backdoor samples $\hat{x}_{i}$ that match the template. Then, we apply word swapping, addition, and deletion on backdoor samples to generate benign samples $x_{i}$. (b) Backdoor Injection: we inject the backdoor via jointly optimizing three losses: $\mathcal{L}_{Benign}$ for minimizing the embedding distance of benign samples between the clean text encoder $F_{\theta}(\cdot)$ and the backdoor text encoder $\hat{F}_{\theta}(\cdot)$, $\mathcal{L}_{Backdoor}$ for aligning the embedding of backdoor samples to target sample, $\mathcal{L}_{KMMDR}$ for providing the gradient information from the U-Net [44] for better aligning cross-attention responses of the diffusion process between backdoor and benign samples.
III-A Overview of Our Methods
@@
Following prior works [25, 27, 26, 34], we consider Text-to-Image (T2I) diffusion models [6] as the threat models for backdoor attacks. In this scenario, the attacker, who has access to the model parameters, injects backdoors into the model and subsequently releases the modified model on the third-party platforms [15]. The defender aims to detect backdoor samples during testing process.
@@
The overview of our method is shown in Fig. 1. Given a T2I diffusion model, the goal is to inject the backdoors into the text encoder $F_{\theta}(\cdot)$ of T2I diffusion models (i.e., CLIP [45]). The backdoor model $\hat{F}_{\theta}(\cdot)$ is expected to output embedding that guides the T2I diffusion model to generate attacker-specified images while mitigating abnormal consistencies. Specifically, our method involves two stages, i.e., data generation and backdoor injection. During the data generation, we first construct a dataset of backdoor samples $\mathbb{\hat{D}}=\{\hat{x}_{i},\hat{y}\}$, where $\hat{x}_{i}$ represents the prompt with the specific syntactic structure $s$ and $\hat{y}$ denotes the attacker-specified target prompt. Benign samples $\mathbb{D}=\{x_{i}\}$ are then generated by applying word-order swapping, words addition, and words deletion transformations to the backdoor samples $\hat{x}_{i}$. For the backdoor injection, the clean model $F_{\theta}(\cdot)$ is fine-tuned on both $\mathbb{\hat{D}}$ and $\mathbb{D}$ to obtain the backdoor model $\hat{F}_{\theta}(\cdot)$. Besides, a regularization loss $\mathcal{L}_{KMMDR}$ based on Kernel Mean Matching Distance (KMMD) is proposed for mitigating the abnormal attention consistency. The overall objective of our method is formulated as follows:
@@
III-B Samples Generation w & w/o Syntactic Trigger
@@
Trigger Syntactic Templates Selection. Based on previous researches [25, 35], we select rare syntactic templates $s$ to generate backdoor samples that effectively distinguish from benign samples. Specifically, we analyze DiffusionDB dataset [46], which contains high-quality prompts from real users. We then leverage the Stanford Parser [47] to parse each prompt and calculate the frequency of various syntactic structures. We choose structures that occur infrequently, such as (DET) (NOUN) (ADP) (DET) (NOUN) (VERB) (ADP) (NOUN), to serve as our templates.
Figure 2: The visualization of cross-attention maps during image generation. Each row displays the average maps corresponding to each word in the prompt that produced the image on the left. (Top) A benign sample. (Middle) A backdoor sample trained by Rickrolling [25]. (Bottom) A backdoor sample with the syntactic trigger trained by our method.
@@
Backdoor Samples Generation. With the selected syntactic templates, we aim to generate a set of backdoor samples $\mathbb{\hat{D}}$ that conform to these syntactic structures. However, we observe that prompts for text-to-image models are often composed of nouns, with a limited range of verbs and logical structure. This makes it challenging to generate human-readable prompts through rule-based methods [35]. To address this, we leverage the off-the-shelf LLM $\mathcal{G}$ [43] to instruct the model to generate sentences with low perplexity that are human-readable, cover varying scenarios, and incorporate varied vocabulary. Thus, we get $\mathbb{\hat{D}}=\mathcal{G}(s)$. Finally, we manually filter out incorrectly generated samples.
Benign Samples Generation. Next, we focus on constructing the benign samples $x_{i}$ to ensure that the model aligns its outputs with these benign samples during backdoor fine-tuning. Instead of directly sampling prompts from DiffusionDB [46], we generate benign samples by applying operations $\mathcal{T}$ such as word-order swapping, words addition, and words deletion to the backdoor samples $\hat{\mathbb{D}}$, i.e., $\mathbb{D}=\mathcal{T}(\hat{\mathbb{D}})$. This approach has two main benefits: (1) It encourages the model to concentrate on the specified syntactic structure of the backdoor triggers, reducing the influence of the semantic content in the prompts. (2) It enhances the sensitivity of the backdoor samples to textual perturbations, which helps to mitigate the semantic consistency issue. We conduct the experiment to show its effectiveness in Section IV-D.
III-C Backdoor Injection
During the backdoor injection process, we utilize a teacher-student paradigm to implant the backdoor, where both models are initialized using a pre-trained CLIP [45]. The student model serves as the poisoned model, while the teacher model is used to align the feature embeddings of samples from $\mathbb{\hat{D}}$ and $\mathbb{D}$ with the student model. It is noted that TwT also involves dual-modal optimization, leveraging gradient from the UNet [44] during the diffusion generation process to jointly optimize the backdoor injection. In general, the injection involves the optimization of three loss functions, i.e., Benign Loss, Backdoor Loss and KMMDR Loss.
@@ III-C Backdoor Injection
where $x,y\in[1,D]$ are the horizontal and vertical coordinates of the attention map, respectively. $\bar{M}$ is the mean of the attention maps across tokens. Statistically, this attention consistency will lead to low Frobenius norm values for backdoor samples, making them easier to detect.
Algorithm 1 TwT Training Procedure
1:trigger template $s$, target prompt $\hat{y}$, off-the-shelf LLM $\mathcal{G}$, prompt operations $\mathcal{T}$, hyperparameters $\gamma$ and $\lambda$, clean model $F_{\theta}(\cdot)$, total number of benign samples $N_{1}$, total number of backdoor samples $N_{2}$, similarity function $d(\cdot,\cdot)$, iterations $K$.
2:Generate Dataset: $\hat{\mathbb{D}}=\mathcal{G}(s)$, $\hat{\mathbb{D}}=\mathcal{T}(\mathbb{D})$;
3:  $\mathbb{D}=\{x_{1},x_{2},\dots,x_{N_{1}}\}$; $\triangleright$ Benign samples
4:  $\mathbb{\hat{D}}=\{\hat{x}_{1},\hat{x}_{2},\dots,\hat{x}_{N_{2}},\hat{y}\}$; $\triangleright$ Backdoor samples
5:Initialize Backdoor Model: $\hat{F_{\theta}}(\cdot)=F_{\theta}(\cdot)$;
6:for $k=1$ to $K$ do
7:     $\mathcal{L}_{Benign}=\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}d(\hat{F}_{\theta}(x_{i}),F_{\theta}(x_{i}))$;
8:     $\mathcal{L}_{Backdoor}=\frac{1}{N_{2}}\sum_{i=1}^{N_{2}}d(\hat{F}_{\theta}(\hat{x}_{i}),F_{\theta}(\hat{y}))$;
9:     for $i=1$ to $N_{1}$ do
10:         Obtain $M_{i}=\{M^{1}_{i},M_{i}^{2},\dots,M_{i}^{L}\}$; $\triangleright$ Attention maps
11:         Compute $P_{i}=\{P_{i}^{1},P_{i}^{2},\dots,P_{i}^{L}\}$;
12:         Compute $C_{i}=\frac{1}{L-1}\sum_{j=0}^{L-1}(P_{i}^{j}-\bar{P})(P_{i}^{j}-\bar{P})^{T}$; $\triangleright$ Prompt length $L$ for sample $x_{i}$
13:     end for
14:     $C=\{C_{1},C_{2},\dots,C_{N_{1}}\}$;
15:     Similarly, $\hat{C}=\{\hat{C}_{1},\hat{C}_{2},\dots,\hat{C}_{N_{2}}\}$;
16:     $\mathcal{L}_{KMMDR}=\text{KMMD}^{2}(C,\hat{C})$ in Eq. (6);
17:     $\mathcal{L}=\mathcal{L}_{Benign}+\gamma\cdot\mathcal{L}_{Backdoor}+\lambda\cdot\mathcal{L}_{KMMDR}$;
18:     Update $\hat{F}_{\theta}(\cdot)$ by $\nabla_{\mathcal{L}}$;
19:end for
20:return $\hat{F}_{\theta}(\cdot)$; $\triangleright$ Backdoor model
@@
Finally, the optimization objective of our method is:
@@
where $\gamma$ and $\lambda$ are parameters that balance the contributions of the Benign Loss, Backdoor Loss and KMMDR Loss. The whole training procedure is illustrated in Algorithm 1.
IV Experiments
In this section, we first introduce the settings of experiments in Section IV-A. Then we demonstrate the effectiveness of our proposed method from Sections IV-B to IV-C and conduct comprehensive ablation study of DAA in Section IV-D. Finally, we discuss the generalization of our method in Section IV-E.
Figure 4: Qualitative results of our method. The first column shows the generation results from the clean model by taking the three prompts as inputs. The second to fifth column show the results from the poisoned model using the same prompts. Each poisoned model is injected with a different backdoor target. We provide more examples in the supplementary material.